OpenStudy (anonymous):

M is the midpoint of JK. The coordinates of J are (6, 3) and the coordinates of M are (-3, 4), find the coordinates of K.

OpenStudy (anonymous):

@Kreshnik .. Help(:

OpenStudy (anonymous):

may I

OpenStudy (anonymous):

Sure .

OpenStudy (anonymous):

\[2 ---> (3-1) & (4-2)\] add the same to point M to get to point K \[x = (3+2) & y = (4+2)\] \[K= (5,6)\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

wel come